OpenStudy (marcelie):

help pleaseeeee Write an equation for a line parallel to f (x) = − 5x − 3 and passing through the point (2, –12).

OpenStudy (anonymous):

Do you know how to plot your graph?

OpenStudy (marcelie):

yes

OpenStudy (lynfran):

hint when two lines are parallel their slope are equal...(they have the same slope) equation of a line ==> y=mx+b when m=slope and b = y-intercept

OpenStudy (marcelie):

is my answer : f(x)= -5x-2 ?

OpenStudy (lynfran):

points (2,-12) m=-5 y=mx+b -12=(-5)(2)+b -12=-10+b -12+10=b -2=b yes y=-5x-2

OpenStudy (marcelie):

Okay. I need help with this one . How do i solve it ? Find the point at which the line f (x) = 2x + 5 intersects the line g(x) = − 3x − 5.

OpenStudy (anonymous):

okay, can you graph both of the lines?

OpenStudy (lynfran):

to find where they intersect we equate the two function and solve for x ... 2x+5=-3x-5

OpenStudy (lynfran):

2x+5=-3x-5 2x+3x=-5-5 5x=-10 x=-2

OpenStudy (lynfran):